# Hurst Index of Gamma-Ray Burst Light Curves and Its Statistical Study

### v1

@@ Abstract
Gamma-ray bursts (GRBs) rank among the most powerful astrophysical phenomena, characterized by complex and highly variable prompt emission light curves that reflect the dynamics of their central engines. In this work, we analyze a sample of 163 long-duration GRBs detected by the Burst and Transient Source Experiment (BATSE), applying detrended fluctuation analysis (DFA) to derive the Hurst index as a quantitative descriptor of temporal correlations in the light curves. We further explore statistical correlations between the Hurst index and 12 other observational parameters through regression and correlation analyses. Our results reveal anti-correlations between the Hurst index and the burst durations (T50, T90), and a negative trend with the low-energy spectral index $\alpha$. We also find that correlations with peak photon flux are strongest at the shorter timescale (64 ms) and systematically weaken at longer timescales (256–1024 ms), indicating that the persistence of temporal correlations is most evident in the rapid variability component of GRB emission. The results offer new perspectives on the temporal structure of the GRB emission and its potential link to the underlying physical mechanisms driving these bursts.
@@ 1 Introduction
Over the years, numerous indicators have been developed to quantify the temporal variability of GRBs. These include pulse decomposition analysis (Norris et al., 1996), variability indices (Fenimore and Ramirez-Ruiz, 2000b), Fourier power density spectra (Beloborodov et al., 2000), autocorrelation functions (Borgonovo, 2004), and minimum variability time scales (MacLachlan et al., 2012). Crucially, some studies have found significant correlations between variability indicators and other properties such as luminosity (Reichart et al., 2001; Guidorzi et al., 2005), suggesting that light curve structure may provide key diagnostics of jet composition and energy dissipation.
@@
Originally introduced by Peng et al. (1994, 1995), the detrended fluctuation analysis (DFA) is a fundamental method to examine the scalar properties in a variety of time series datasets. Later, Kantelhardt et al. (2002) expanded DFA to analyze multifractal processes, resulting in the multifractal detrended fluctuation analysis (MFDFA) technique. A generalized Hurst index was defined to quantify the long-range dependence of a time series, which was initially introduced by Hurst (1951) in the study of the long-term storage capacity of reservoirs. This method effectively addresses correlation issues in time-series data and is applicable to both discrete and continuous stochastic processes. It has been used successfully in a wide range of applications spanning several fields, such as music (Jafari et al., 2007), heartbeat dynamics (Ivanov et al., 1999), electroencephalogram sleep data (Pavlov et al., 2020b, 2021), arterial pressure (Pavlov et al., 2020a), cosmic microwave radiation (Movahed et al., 2011, 2013), atmospheric turbulence effects on stellar images (Zunino et al., 2014), sunspot fluctuations (Sadegh Movahed et al., 2006; Hu et al., 2009), solar flares (Lee et al., 2020), and gravitational wave detection (Eghdami et al., 2018). In this context, DFA and its extension to MFDFA offer powerful tools to characterize temporal correlations and long-range memory in GRB light curves. Unlike traditional variability measures, DFA quantifies the scaling behavior and persistence of non‑stationary time series, making it particularly suited to GRB data (Peng et al., 1994; Kantelhardt et al., 2002; Zunino et al., 2014).
@@
Since its launch in 1991 onboard NASA’s Compton Gamma Ray Observatory, the Burst and Transient Source Experiment (BATSE) has been instrumental in shaping current understanding of GRBs (Band et al., 1993). One of BATSE’s most impactful achievements is its systematic collection of high-resolution light curves in four well-defined energy bands: 20-50 keV, 50-100 keV, 100-300 keV and above 300 keV, which has facilitated detailed multiband temporal and spectral investigations of GRB prompt emission. Using this extensive data set, Hakkila (2021) found that most GRB pulses can be well characterized by a smooth, single-peaked emission profile accompanied by a temporally symmetric residual component, shedding new light on the structure and physical origin of GRB emission mechanisms. We also chose the BATSE data for our analysis because of the high sensitivity.
@@
In this paper, we analyze the detrended fluctuations in the light curves of 163 long GRBs detected by BATSE, obtain a characteristic index (Hurst index) consisting of the light curves as a time series, and perform regression and correlation analyses between this index and 12 different physical parameters that have been observed for these long GRBs. The paper is arranged as follows. In Section 2, we present our data sample and describe the 12 physical parameters for which we performed the correlation analysis. The detrended fluctuation analysis model and its parameter selection and statistical analysis are presented in Section 3. The results and conclusion are reported in Sections 4 and 5, respectively.
@@ 2.1 Sample selection
Our analysis is based on a carefully selected sample of 163 long-duration GRBs detected by BATSE, corresponding to trigger numbers ranging from 105 to 1999111https://heasarc.gsfc.nasa.gov/FTP/compton/data/batse/ascii_data/64ms/, which is the intersection of BATSE Gamma Ray Burst Lightcurve Image Archive [^2] https://gammaray.nsstc.nasa.gov/batse/grb/lightcurve/ and CGRO/BATSE 4B Catalog [^3] https://heasarc.gsfc.nasa.gov/W3Browse/cgro/batse4b.html. The sample selection criteria also included: complete determinations of the spectral parameters from band function fittings and availability of all 12 other physical parameters under investigation from Wang et al. (2020). The 12 selected parameters are shown in Table 1. The light curves are rebinned to a uniform time resolution of 1024 ms to standardize temporal sampling across all gamma-ray bursts. The time interval for each burst is selected from the trigger time up to $T_{90}$. This standardization ensures that the analyzed data encompass the entire prompt emission phase, guaranteeing that the scaling properties derived from detrended fluctuation analysis are representative of the burst’s intrinsic variability during its most active phase and enabling meaningful comparison of Hurst indices for robust statistical evaluation. Fig. 1 presents the light curve of GRB 920110A as an example, one of the events included in our sample, over the time span corresponding to its $T_{90}$ duration. The A provides detailed information on all gamma-ray bursts included in our sample, along with the corresponding physical parameters from Wang et al. (2020) analyzed in this study.
@@ 2.2 Physical Parameters
We analyze 12 key physical parameters that characterize different aspects of GRB emission. These parameters can be broadly categorized into three groups: (1) spectral parameters describing the energy distribution of photons, (2) temporal parameters characterizing the burst duration and variability, and (3) flux-related parameters measuring the burst intensity. Table 1 summarizes these parameters with their units and physical interpretations. The spectral parameters $\alpha$ and $\beta$ represent the low- and high-energy photon indices from the Band function fits, while $E_{peak}$ corresponds to the peak energy in the $\nu F_{\nu}$ spectrum. The temporal parameters $T_{50}$ and $T_{90}$ describe the durations of the bursts that contain total fluence 50% and 90%, respectively, and V quantifies the degree of fluctuations in the light curve. The flux-related parameters include fluence ($F_{g}$), peak flux ($F_{pk}$), and photon flux ($P_{pk}$) measured in different time bins. The hardness ratio (HR) provides information about the spectral hardness of the burst. All data were taken from Wang et al. (2020), while they were gathered from references therein.
@@
$F_{g}$ $10^{-9}$ J m-2 Fluence in the 20-2000 keV energy band
$F_{pk}$ $10^{-9}$ J m-2 s-1 Peak flux in the 1 s time bin in the restframe 1-10000 keV energy band
@@
$T_{50}$ s Burst duration containing 25%-75% of the total fluence
$T_{90}$ s Burst duration containing 5%-95% of the total fluence
@@
which also corresponds to variablity_1 in Wang et al. (2020)
@@ 3 Methods
3.1 Detrended Fluctuatioin Analysis
The DFA method was originally developed by Peng et al. (1994) and was later extended to multifractal analysis by Kantelhardt et al. (2002). Following Peng et al. (1994) and Kantelhardt et al. (2002), the DFA procedure consists of the following steps:
@@ 3.1 Detrended Fluctuatioin Analysis
1. For a given light curve represented as a time series $x(i)$ of length $N$, we first construct the integrated series:
$y(k)=\sum_{i=1}^{k}[x(i)-\langle x\rangle],$ (1)
where $\langle x\rangle$ is the mean value of the time series.
2. The integrated series is divided into non-overlapping segments of equal length $s$. In each segment $\nu$, we fit a polynomial trend $y_{\nu}(k)$ of the order $n$ (typically linear or quadratic) to remove local trends.
3. The detrended fluctuation function $F(\nu,s)$ for each segment is calculated as:
$F^{2}(\nu,s)=\frac{1}{s}\sum_{k=(\nu-1)s+1}^{\nu s}[y(k)-y_{\nu}(k)]^{2}.$ (2)
4. The overall fluctuation function $F(s)$ is obtained by averaging across all segments:
$F(s)=\left[\frac{1}{2N_{s}}\sum_{\nu=1}^{2N_{s}}F^{2}(\nu,s)\right]^{1/2},$ (3)
where $N_{s}=\text{int}(N/s)$.
5. The process is repeated for different time scales $s$ to examine the scaling behavior. For fractal processes, $F(s)$ follows a power law:
$F(s)\sim s^{H},$ (4)
where $H$ is the Hurst index, which serves as our primary measure of temporal correlation. The Hurst index ranges between 0 and 1, with $H=0.5$ indicating uncorrelated (white noise) behavior, $H>0.5$ indicating persistent long-range correlations, and $H<0.5$ indicating anti-persistent correlations.
Following Kantelhardt et al. (2002) and Rydin Gorjão et al. (2022), for multifractal analysis, we generalize the fluctuation function to the $q$-th order:
$\mathcal{F}(q,s)=\left(\frac{1}{2N_{s}}\sum_{\nu=1}^{2N_{s}}\left[F^{2}(\nu,s)\right]^{q/2}\right)^{1/q}.$ (5)
For $q=0$, we use logarithmic averaging:
$\mathcal{F}(0,s)=\exp\left(\frac{1}{4N_{s}}\sum_{\nu=1}^{2N_{s}}\ln\left[F^{2}(\nu,s)\right]\right).$ (6)
The Hurst index $H$ is particularly relevant for GRB light curves, as it provides a quantitative measure of “memory” in the emission process. A high Hurst index suggests that the central engine activity exhibits long-term correlations, possibly indicative of sustained accretion or magnetic processes, while a low Hurst index might imply more stochastic behavior dominated by short-timescale phenomena.
3.2 Implementation and Parameter Optimization
Our analysis was implemented using the MFDFA Python package (Rydin Gorjão et al., 2022), which provides efficient computation of both standard DFA and multifractal extensions [^4] https://github.com/LRydin/MFDFA. The key parameters in our implementation include:
1. Scale selection: We used logarithmically spaced segment sizes (scales) ranging from a minimum of 3 data points ($0.192$ s) to a maximum of one-tenth the length of each light curve, with 200 scale points. This range ensures adequate statistics while avoiding both undersampling at small scales and excessive averaging at large scales.
2. Detrending: We used first-order polynomial detrending (linear detrending, order=1) to remove local trends while preserving the intrinsic variability of the light curves. This choice represents the standard approach in DFA implementations for astrophysical time-series analysis.
3. Fluctuation order: For the monofractal analysis presented here, we used $q=2$, which corresponds to the conventional DFA implementation and provides the Hurst index as a measure of long-range correlation.
4. Data processing: Each gamma-ray burst light curve was analyzed within the astrophysical relevant time window from trigger time ($t=0$) to duration $T_{90}$ (the time interval containing 90% of the total burst emission). This ensures that the analysis focuses on the most significant emission phase while excluding background-dominated regions. The counts were analyzed in both their total (integrated across all channels) and individual BATSE energy channels (channels 1-4) to investigate potential energy-dependent correlations.
5. Hurst index calculation: The Hurst index $H$ was derived from the slope of the fluctuation function $F(s)$ versus the size of the segment $s$ in log-log space using ordinary least squares regression. Following the standard DFA convention for nonstationary time series (Kantelhardt et al., 2002), $H=\alpha-1$, where $\alpha$ is the fitted slope of $\log(F(s))$ versus $\log(s)$.
6. Uncertainty estimation: The standard errors of the Hurst indices were estimated using the ordinary least-squares regression process. For cases where regression-based error estimation failed (typically due to very short data segments), we used resampling techniques to ensure reliable uncertainty quantification.
Following the methodology described above, we computed the Hurst indices for all four energy channels of 163 GRBs from the BATSE catalog. In this work, we choose the combined light curves of all four channels. It should be noted that we fixed the parameter $q$ to 2, which means that we performed DFA rather than MFDFA. However, the whole method described above could be used as a general treatment for further research in the future. Table 2 presents a representative sample of these results, showing the trigger number, the name of the GRB, and the derived Hurst index ($H$) for all channels. The complete Hurst index list for our sample is presented in A.
Trigger number GRB name Hurst index
1982 GRB 921008 $0.73\pm 0.07$
1989 GRB 921015 $0.58\pm 0.04$
1991 GRB 921017 $0.54\pm 0.07$
Table 2: Representative sample of results.

### v4

@@ Abstract
Gamma-ray bursts (GRBs) rank among the most powerful astrophysical phenomena, characterized by complex and highly variable prompt emission light curves that reflect the dynamics of their central engines. In this work, we analyze a sample of 163 long-duration GRBs detected by the Burst and Transient Source Experiment (BATSE), applying detrended fluctuation analysis (DFA) to derive the Hurst index as a quantitative descriptor of temporal correlations in the light curves. We further explore statistical correlations between the Hurst index and 12 other observational parameters through regression and correlation analyses. Our results reveal anti-correlations between the Hurst index and the burst durations ($T_{50}$, $T_{90}$), and moderate positive correlations with peak photon flux proxies (Ppk1–Ppk3). By contrast, the standard spectral parameters (including the low-energy index $\alpha$) show no evidence for a linear dependence on the Hurst index in our sample. We do not find a clear monotonic weakening of the correlation strength from 64 ms to 1024 ms peak-flux measures; rather, the correlation coefficients for Ppk1–Ppk3 are comparable within uncertainties. The results offer new perspectives on the temporal structure of the GRB emission and its potential link to the underlying physical mechanisms driving these bursts.
@@ 1 Introduction
Over the years, numerous indicators have been developed to quantify the temporal variability of GRBs. These include pulse decomposition analysis (Norris et al., 1996), variability indices (Fenimore and Ramirez-Ruiz, 2000b), Fourier power density spectra (Beloborodov et al., 2000), autocorrelation functions (Borgonovo, 2004), and minimum variability time scales (MacLachlan et al., 2012). Crucially, some studies have found significant correlations between variability indicators and other properties such as luminosity (Reichart et al., 2001; Guidorzi et al., 2005), suggesting that the light curve structure may provide key diagnostics of jet composition and energy dissipation.
@@
Originally introduced by Peng et al. (1994, 1995), detrended fluctuation analysis (DFA) is a fundamental method for examining the scalar properties in a variety of time series datasets. Later, Kantelhardt et al. (2002) expanded DFA to analyze multifractal processes, resulting in the multifractal detrended fluctuation analysis (MFDFA) technique. A generalized Hurst index was defined to quantify the long-range dependence of a time series, initially introduced by Hurst (1951) in the study of the long-term storage capacity of reservoirs. This method effectively addresses correlation issues in time-series data and is applicable to both discrete and continuous stochastic processes. It has been used successfully in a wide range of applications spanning several fields, such as music (Jafari et al., 2007), heartbeat dynamics (Ivanov et al., 1999), electroencephalogram sleep data (Pavlov et al., 2020b, 2021), arterial pressure (Pavlov et al., 2020a), cosmic microwave radiation (Movahed et al., 2011, 2013), atmospheric turbulence effects on stellar images (Zunino et al., 2014), sunspot fluctuations (Sadegh Movahed et al., 2006; Hu et al., 2009), solar flares (Lee et al., 2020), gravitational wave detection (Eghdami et al., 2018), quasiperiodic oscillation searching (Tarnopolski and Marchenko, 2021), blazars (Tarnopolski et al., 2020), and fast radio bursts (Wang et al., 2023). In this context, DFA and its extension to MFDFA offer powerful tools to characterize temporal correlations and long-range memory in GRB light curves. Unlike traditional variability measures, DFA quantifies the scaling behavior and persistence of non‑stationary time series, making it particularly suited to GRB data (Peng et al., 1994; Kantelhardt et al., 2002; Zunino et al., 2014).
@@
Since its launch in 1991 onboard NASA’s Compton Gamma Ray Observatory, the Burst and Transient Source Experiment (BATSE) has been instrumental in shaping our current understanding of GRBs (Band et al., 1993). One of BATSE’s most impactful achievements is its systematic collection of high-resolution light curves in four well-defined energy bands: 20-50 keV, 50-100 keV, 100-300 keV, and above 300 keV, which has facilitated detailed multiband temporal and spectral investigations of GRB prompt emission. Using this extensive data set, Hakkila (2021) found that most GRB pulses can be well characterized by a smooth, single-peaked emission profile accompanied by a temporally symmetric residual component, shedding new light on the structure and physical origin of GRB emission mechanisms. We also chose the BATSE data for our analysis because of its high sensitivity.
@@
In this paper, we analyze the detrended fluctuations in the light curves of 163 long GRBs detected by BATSE, derive a characteristic index (the Hurst index) from each GRB light curve treated as a time series, and perform regression and correlation analyses between this index and 12 different physical parameters that have been observed for these long GRBs. The paper is arranged as follows. In Section 2, we present our data sample and describe the 12 physical parameters for which we performed the correlation analysis. The detrended fluctuation analysis model, along with its parameter selection and statistical analysis, is presented in Section 3. The results and conclusions are reported in Sections 4 and 5, respectively.
@@ 2.1 Sample Selection
Our analysis is based on a carefully selected sample of 163 long-duration GRBs detected by BATSE, corresponding to trigger numbers ranging from 107 to 1997111https://heasarc.gsfc.nasa.gov/FTP/compton/data/batse/ascii_data/64ms/, which is the intersection of the BATSE Gamma Ray Burst Lightcurve Image Archive [^2] https://gammaray.nsstc.nasa.gov/batse/grb/lightcurve/ and the CGRO/BATSE 4B Catalog [^3] https://heasarc.gsfc.nasa.gov/W3Browse/cgro/batse4b.html. The sample selection criteria also included: complete determinations of the spectral parameters from band function fittings and availability of all 12 other physical parameters under investigation from Wang et al. (2020). The 12 selected parameters are shown in Table 1. The DISCSC light curves are originally provided at 64 ms resolution. In this work, we rebin the DISCSC data to a uniform time resolution of 1024 ms by summing 16 consecutive 64 ms bins, so that all bursts are analyzed on the same temporal grid. This uniform binning is adopted because DFA results are sensitive to the sampling resolution; fixing the bin width across the sample avoids introducing an additional burst-dependent scale that could bias the distribution of Hurst indices and compromise the comparability required for our statistical study.
This choice also imposes a practical requirement on the minimum number of data points available for each burst. Starting from an initially larger candidate set, we excluded 36 GRBs with $T_{90}<8.6$ s, for which the rebinned light curves would contain too few 1024 ms bins within the prompt-emission window to yield a meaningful DFA scaling fit. After this filtering, 163 long GRBs remain and constitute our final sample. We note that relatively short bursts (e.g., those with $T_{90}\lesssim 30$ s) are still included as long as they satisfy the above minimum-length requirement under the uniform 1024 ms binning.
The time interval for each burst is selected from the trigger time up to $T_{90}$. This standardization provides a uniform and reproducible analysis window, guaranteeing that the scaling properties derived from detrended fluctuation analysis are representative of the burst’s intrinsic variability during its most active phase and enabling a meaningful comparison of Hurst indices for robust statistical evaluation. Fig. 1 presents the light curve of GRB 920110A as an example, one of the events included in our sample, over the time span corresponding to its $T_{90}$ duration. A (LABEL:tab:list_HI) lists the Hurst indices for all GRBs in our sample; the corresponding physical parameters used in the correlation analysis are taken from Wang et al. (2020). We emphasize that in this work $T_{90}$ is used as a practical duration parameter to define an analysis window that covers the vast majority of the prompt-emission variability, rather than as a strict physical gate that begins at the canonical $t_{\rm start}$ of the $T_{90}$ definition. In many BATSE long GRBs, the light curve shows a clear rise and non-negligible variability immediately after trigger ($t\approx 0$), and excluding the early segment prior to the onset of the cataloged $T_{90}$ interval would remove part of the burst structure that can contribute to the scaling behavior measured by DFA. Therefore, we adopt the interval $t\in[0,\,T_{90}]$ for all bursts to ensure methodological uniformity across the sample and to avoid selectively discarding early-time variability. The purpose of our study is to obtain a standard set of DFA-based scaling measures for statistical comparison across the sample. In Fig. 1, the dashed line marks the endpoint of the adopted analysis window at $t=T_{90}$ (measured from the trigger). The caption explicitly clarifies that $T_{90}$ is a duration parameter, and that the vertical marker indicates the end of the $0,\,T_{90}]$ window rather than the $T_{5},\,T_{95}]$ window. For five triggers (No. 107, 110, 114, 211, 351) whose publicly available DISCSC records terminate before $t=T_{90}$, we perform DFA on the actually available window $0,\,t_{\rm end}^{\rm data}]$ and treat the cataloged $T_{90}$ as a reference duration parameter.
@@ 2.2 Physical Parameters
We analyze 12 key physical parameters that characterize different aspects of GRB emission. These parameters can be broadly categorized into three groups: (1) spectral parameters describing the energy distribution of photons, (2) temporal parameters characterizing the burst duration and variability, and (3) flux-related parameters measuring the burst intensity. Table 1 summarizes these parameters with their units and physical interpretations. The spectral parameters $\alpha$ and $\beta$ represent the low- and high-energy photon indices from the Band function fits, while $E_{peak}$ corresponds to the peak energy in the $\nu F_{\nu}$ spectrum. The temporal parameters $T_{50}$ and $T_{90}$ describe the durations of the bursts that contain total counts of 50% and 90%, respectively, and V quantifies the degree of fluctuations in the light curve. The flux-related parameters include fluence ($F_{g}$), peak flux ($F_{pk}$), and photon flux ($P_{pk}$) measured in different time bins. The hardness ratio (HR) provides information about the spectral hardness of the burst. All data were taken from Wang et al. (2020), while they were gathered from the references therein.
@@
$F_{g}$ $10^{-6}$ erg cm-2 Fluence in the 20-2000 keV energy band
$F_{pk}$ $10^{-6}$ erg cm-2 s-1 Peak flux with 1 s time bin in the observer’s frame 1-10000 keV energy band
@@
$T_{50}$ s Burst duration between 25% and 75% of the cumulative counts (following Wang et al. (2020))
$T_{90}$ s Burst duration between 5% and 95% of the cumulative counts (following Wang et al. (2020))
@@
which also corresponds to variability1 in Wang et al. (2020)
@@ 3 Methods
3.1 Detrended Fluctuation Analysis (DFA)
Detrended fluctuation analysis (DFA) was originally proposed by Peng et al. (1994) and later generalized to multifractal detrended fluctuation analysis (MF-DFA) by Kantelhardt et al. (2002). In this work, we focus on the monofractal case (DFA) by fixing $q=2$ in the MF-DFA formalism.
@@ 3.1 Detrended Fluctuation Analysis (DFA)
Given a GRB light curve represented by a discrete time series $x(i)$ of length $N$, DFA proceeds as follows:
(1) Profile construction. We construct the cumulative (integrated) profile
$y(l)=\sum_{i=1}^{l}\left[x(i)-\langle x\rangle\right],$ (1)
where $\langle x\rangle$ is the mean of the time series. Here $x(i)$ is the binned photon counts (counts per bin) in the $i$-th time bin of width $\Delta t=1.024~\mathrm{s}$ (after summing the four DISCSC channels).
(2) Segmentation and detrending. For a chosen scale (window size) $s$, the profile $y(l)$ is divided into $N_{s}=\lfloor N/s\rfloor$ non-overlapping segments of length $s$. To reduce boundary effects and to use the entire time series, we perform the same segmentation starting from the end of the profile, yielding a total of $2N_{s}$ segments (Kantelhardt et al., 2002). In each segment $\nu$, we fit an $n$-th order polynomial trend $y_{\nu}^{(n)}(l)$ (in this work $n=1$) via least squares and subtract it from the profile.
(3) Local variance. For each segment $\nu$, the detrended variance is computed as
$F^{2}(\nu,s)=\frac{1}{s}\sum_{k=1}^{s}\left[y_{\nu}(l)-y_{\nu}^{(n)}(l)\right]^{2},$ (2)
where $y_{\nu}(l)$ denotes the profile values restricted to segment $\nu$.
(4) Fluctuation function. The fluctuation function at scale $s$ is obtained by averaging over all $2N_{s}$ segments:
$F(s)=\left[\frac{1}{2N_{s}}\sum_{\nu=1}^{2N_{s}}F^{2}(\nu,s)\right]^{1/2}.$ (3)
(5) Scaling and slope estimation. Repeating the above steps for a set of scales $\{s\}$, DFA predicts a power-law scaling
$F(s)\propto s^{k},$ (4)
where $k$ is the fitted slope in the $\ln F(s)$$\ln s$ plane. We estimate $k$ using ordinary least squares (OLS) regression over all available scales used in the computation (see Section 3.2). To avoid confusion with the low-energy spectral index $\alpha$ of the Band function, we denote the DFA scaling slope by $k$ throughout this paper.
Following the convention implemented in the MFDFA package (Rydin Gorjão et al., 2022), we report the Hurst index for the light-curve series as
H=k1.H=k-1. (5)
Here $H$ is an empirical DFA scaling index defined by Eq. (5). For finite-length and strongly nonstationary burst light curves, the fitted scaling exponent may fall outside the canonical range expected for idealized fGn/fBm, and such cases should be interpreted with caution.
3.2 Implementation and analysis configuration
Time resolution and rebinning
The original BATSE light-curve data used in this work have a native time bin width of 64 ms. Prior to the DFA computation, we rebin each burst to a uniform time resolution of 1024 ms (1.024 s) by summing consecutive 16 bins. For a discrete count series $x_{64}(i)$ sampled at 64 ms, the rebinned series is
$x_{1024}(j)=\sum_{m=1}^{16}x_{64}\left(16(j-1)+m\right)\quad j=1,2,\dots,N_{1024}.$ (6)
All Hurst indices reported in this paper are derived from the rebinned 1024 ms light curves; the 64 ms data are used only as the input for rebinning. We note that BATSE monitors triggering on 64 ms, 256 ms, and 1024 ms timescales; this triggering timescale does not imply that our DFA is performed on 64 ms bins.
Time window definition
[…]
Scale selection and detrending order
We use logarithmically spaced scales from $s_{\min}=3$ bins to $s_{\max}=\lfloor N/10\rfloor$ bins, with 200 candidate scale points before integer rounding and deduplication. With $\Delta t=1.024$ s, the minimum scale corresponds to $s_{\min}\Delta t=3.072$ s. We adopt first-order polynomial detrending in each segment. For short rebinned series where the nominal upper scale $s_{\max}=\lfloor N/10\rfloor$ falls below $s_{\min}=3$ bins, we retain the burst by fitting over the available integer scales after rounding and deduplication.
Hurst index and uncertainty
For each GRB light curve, we evaluate the fluctuation function $F(s)$ on a set of logarithmically spaced window sizes $s$ and determine the scaling exponent by fitting the linear model
$\ln F(s)=k\,\ln s+b$ (7)
using an unweighted ordinary least-squares (OLS) regression over all adopted scales. We then report the Hurst index as H=k1H=k-1.
When the regression is well defined, the statistical uncertainty of the fitted slope $k$ is quantified by its OLS standard error. Specifically, let $n$ be the number of scale points used in the fit, $y_{i}=\ln F(s_{i})$, and $\hat{y}_{i}$ the fitted values. The residual sum of squares is
$\mathrm{RSS}=\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2},$ (8)
and the mean squared error (i.e., the unbiased estimator of the residual variance) is
$\mathrm{MSE}=\frac{\mathrm{RSS}}{n-2},$ (9)
where n2n-2 is the number of degrees of freedom for a linear fit (slope and intercept). With the design matrix $X=[\ln s_{i},\,1]$, the standard error of the slope is
$\mathrm{SE}(k)=\sqrt{\mathrm{MSE}\,\left[(X^{\mathsf{T}}X)^{-1}\right]_{11}}.$ (10)
Because $H$ differs from $k$ by a constant offset, the propagated uncertainty is simply $\mathrm{SE}(H)=\mathrm{SE}(k)$.
In a small number of cases (typically very short time series), the regression-based uncertainty may be ill-conditioned or numerically unstable. For these bursts, we estimate the uncertainty from repeated resampling of the time series (bootstrap or subsampling, depending on data length) and record the adopted uncertainty-estimation method for each GRB.
Worked example
To make the derivation of $H$ explicit, Figure 2 shows the $\ln F(s)$$\ln s$ scaling relation and the OLS fit for the GRB in Figure 1 (GRB 920110A). The corresponding $(s,F(s))$ data used for the fit are exported from our pipeline to enable full reproducibility.
Figure 2: $\ln F(s)-\ln s$ Scaling Relation and the OLS fit for GRB 920110A